# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v2

@@ 5.2 Global optimality of Lagrangian TMA
As shown below, Lagrangian TMA demonstrably solves the maximisation problem in the robust objective (see Eq. 4) by performing mirror ascent on a suitable PTK. Note that the analysis follows closely that of Wang et al., (2024), where the Lagrangian and the weighted Bregman divergence are explicitly separated in this subsection.
The first step of our analysis of Lagrangian TMA (LTMA) is to derive the gradient of the transition dynamics.
@@
The result follows from Theorem 5.7 in Wang et al., (2024) by considering the Lagrangian as a value function (see Eq. 5). ∎
Note that the Lagrangian adversarial policy gradient in Bossens, (2024) follows similar reasoning but there are two differences in the setting. First, due to the formulation of costs as $c_{0}(s,a,s^{\prime})$ rather than $c_{0}(s,a)$, the state-action value of every time step appears rather than the value of every next time step. Second, we show the proof for a discounted Lagrangian of a potentially infinite trajectory rather than an undiscounted $T$-step trajectory. Third, in the following, we apply it not only to the traditional Lagrangian but also to the augmented Lagrangian and the augmented regularised Lagrangian.
@@
Following Theorem 5.5 in Wang et al., (2024), it follows for any value function with cost in $[0,1]$ that
@@
Reformulating the Lagrangian as a value function (see Eq. 5), considering the bounds on the absolute value of the constraint-cost Eq. 10, and observing that the bounds on the (un-)augmented Lagrangian lie in $F_{\lambda},F_{\lambda}]$ (see Eq. 11 and  12), the result can be scaled to obtain
@@
$\mathbf{V}_{\pi,p^{*}}(\rho;\lambda)-\mathbf{V}_{\pi,p^{t}}(\rho;\lambda)$ $\leq\frac{2F_{\lambda}}{t}\left(\frac{M+1}{(1-\gamma)^{2}}\right)$
$\leq\epsilon^{\prime}\,.$
@@ 5.3 Analysis of Approximate TMA
Definition 2 (Action next-state value function (Definition 3.2 in Li et al., (2023))).
@@
It can also be written as $G_{\pi,p}(s,a,s^{\prime})=c(s,a,s^{\prime})+\gamma V_{\pi,p}(s^{\prime})$ (Wang et al.,, 2024).
@@
The proof follows directly by applying Lemma 5.1 of Wang et al., (2024) to the approximate value function. ∎
@@
$p^{t+1}(\cdot|s,a)=\mathrm{arg}\max_{p\in\Delta(\mathcal{S})}\eta_{p}(t)\frac{1}{1-\gamma}d_{\rho}^{\pi,p^{t}}(s,a)\hat{G}_{\pi,p^{t}}(s,a,\cdot)-\alpha_{p}B_{d_{\rho}^{\pi,p^{t}}}(p,p^{t})\quad\text{for all }\mathcal{S}\times\mathcal{A}\,.$ (26)
As a consequence of this update rule, the value difference in $p^{t+1}$ and $p^{t}$ can be bounded. In particular, the ascent property of TMA (Lemma 5.4 of Wang et al., (2024)) is a key result that can be used to construct the proof for the approximate TMA algorithm. The property, which is given below, notes that conditioning on the state or state-action, a pushback property can be achieved, and this can be used to show that (under perfect gradient information) TMA updates always result in increasing value.
Lemma 4 (Ascent property of TMA (Lemma 5.4 of Wang et al., (2024))).
@@
$\langle G_{\pi,p}(s,\cdot,\cdot),p-p^{t+1}(\cdot|s,\cdot)\rangle+B(p^{t+1}(\cdot|s,\cdot),p^{t}(\cdot|s,\cdot))\leq B(p,p^{t}(\cdot|s,\cdot))-B(p,p^{t+1}(\cdot|s,\cdot))\,$, (27)
@@
$V_{\pi,p^{t}}(\rho)-V_{\pi,p^{t+1}}(\rho)\leq\frac{2C\epsilon^{\prime}}{1-\gamma}\,.$ (30)
@@
$\leq\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\left\lVert G_{\pi,p^{t}}(s,a,\cdot)-\hat{G}_{\pi,p^{t}}(s,a,\cdot)\right\rVert_{\infty}\left(\left\lVert p^{t}(\cdot|s,a)\right\rVert_{1}+\left\lVert p^{t+1}(\cdot|s,a)\right\rVert_{1}\right)$ (triangle inequality)
$\leq\frac{2C\epsilon^{\prime}}{1-\gamma}\,.$ (Eq. 29)
@@
Let $V:\mathcal{S}\to\mathbb{R}$ be a value function with instantaneous cost bounded in [C,C][-C,C], fix $\pi\in\Pi$, and let $p^{*}\in\mathrm{arg}\min_{p\in\mathcal{P}}V_{\pi,p}(\rho)$. Suppose we have an estimate $\hat{G}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ of the action next-state value function which is $\epsilon^{\prime}$-precise as in Eq. 29, and the step sizes follow $\eta_{p}(0)\geq\alpha_{p}(0)\frac{\gamma}{C(1-\gamma)}B_{d_{\rho}^{\pi,p^{0}}}(p^{*},p^{0})$, $\eta_{p}(t)=\eta_{p}(t-1)/\gamma$ and $\alpha_{p}(t)=\alpha_{p}(0)>0$ for all $t\geq 1$. It follows that approximate TMA (Algorithm 1) yields a regret bounded by
@@
12:   $p^{t+1}(\cdot|s,a)=\mathrm{arg}\max_{p\in\Delta(\mathcal{S})}\eta_{p}(t)\left\langle p,\hat{G}_{\pi,p^{t}}(s,a,\cdot)\right\rangle-\alpha_{p}B_{d_{\rho}^{\pi,p^{t}}}(p,p^{t})$ for all $\mathcal{S}\times\mathcal{A}$.
@@ 5.4 Analysis of Robust Sample-based PMD-PD
Having shown results Lagrangian TMA in oracle-based and sample-based settings, we now turn to proving the sample complexity of Robust Sample-based PMD-PD (Algorithm 2). The algorithm modifies Sample-based PMD-PD (Liu et al., 2021b, ) by accounting for the mismatch between the nominal transition kernel and the optimal transition kernel in the min-max problem in Eq. 4. The complexity analysis below is slightly simplified by dropping purely problem-dependent constants $\gamma$, $\eta$, $\zeta$ from the big-O statements.
@@
As sample-based algorithms work with approximate value functions, Lemma 9 describes the conditions under which one can obtain an $\epsilon$-approximation of the cost functions and the regularised augmented Lagrangian. The same settings as sample-based PMD in Lemma 15 of Liu et al., 2021b transfer to the Robust Sample-based PMD algorithm. We rephrase the lemma to include updated transition dynamics.
Lemma 9 (Value function approximation, Lemma 15 of Liu et al., 2021b rephrased).
@@
Lemma 10 (Bound on regularised augmented Lagrangian under approximate entropy-regularised NPG, Lemma 18 in Liu et al., 2021b ).
@@
The proof makes use of Lemma 19. In particular, it notes that $K=\Theta(1/\epsilon)$ and the value difference to $\pi_{k}^{*}$ is bounded by $1/K$ after $t_{k}$ steps. After applying the pushback property (Lemma 16) to $\pi$, $\pi_{k}$, and $\pi_{k}^{*}$, and derives an upper bound on the Bregman divergence w.r.t. $\pi_{k}^{*}$ based on $B_{d_{\rho}^{\pi,p_{k}}}(\pi,\pi_{k+1})+\left\lVert\log(\pi_{k}^{*})-\log(\pi_{k}^{t+1})\right\rVert_{\infty}$, where the second term is $\Theta(1/K)$. We refer to Lemma 18 of Liu et al., 2021b for the detailed proof. ∎
@@
Choose parameter settings according to Lemma 9, let $\alpha=\frac{2\gamma^{2}m\eta_{\lambda}}{(1-\gamma)^{3}}$, $\eta_{\lambda}=1$, and $\eta=\frac{1-\gamma}{\alpha}$. Moreover, let $\pi^{*}\in\mathrm{arg}\inf_{\pi\in\Pi}\Phi(\pi)$, let $\epsilon_{k}^{\prime}>0$ be the upper bound on the LTMA error tolerance at macro-iteration $k$ such that $\epsilon_{k}^{\prime}=\Theta(\epsilon)$, and let $\mathcal{P}$ be an $\ell_{q}$ uncertainty set as in the precondition of Lemma 23. Then within a total number of queries to the generative model equal to $T=\tilde{\mathcal{O}}(\epsilon^{-3})$, Robust Sample-based PMD-PD provides the following guarantees:
(a) regret upper bound:
@@
(b) constraint-cost upper bound:
@@
c) robust Lagrangian regret upper bound under LTMA oracle: under the conditions of Lemma 2, with a number of LTMA iterations $t_{k}^{\prime}=\Theta\left(\frac{F_{\lambda_{k}}M}{\epsilon_{k}^{\prime}(1-\gamma}\right)$, we have
@@
d) robust Lagrangian regret upper bound under approximate LTMA: under the conditions of Lemma 7 as applied to the Lagrangian at each macro-iteration, with $C=F_{\lambda}$, we have
@@
(a) The proof uses Lemma 10 and Lemma 14 to obtain a telescoping sum similar to the analysis of Theorem 3 in Liu et al., 2021a . To account for transition kernel changes, it then upper bounds the cumulative regret difference using Lemma 23) and the Bregman divergence terms in the RHS of Eq. 36 using Lemma 24. This then results in a telescoping sum.
@@
Sample complexity: one may then apply Algorithm 2, plug in the settings from Lemma 9 and Lemma 7 to obtain

### v4

@@ 5.2 Global optimality of Lagrangian TMA
As shown below, Lagrangian TMA (LTMA) demonstrably solves the maximisation problem in the robust objective (see Eq. 4) by performing mirror ascent on a suitable PTK. Note that the analysis follows closely that of Wang et al. (2024), where the Lagrangian and the weighted Bregman divergence are explicitly separated in this subsection.
The first step of our analysis of LTMA is to derive the gradient of the transition dynamics.
@@
The result follows from Theorem 5.7 in Wang et al. (2024) by considering the Lagrangian as a value function (see Eq. 5). ∎
Note that the Lagrangian adversarial policy gradient in Bossens (2024) follows similar reasoning but there are two differences in the setting. First, due to the formulation of costs as $c_{0}(s,a,s^{\prime})$ rather than $c_{0}(s,a)$, the state-action value of every time step appears rather than the value of every next time step. Second, we show the proof for a discounted Lagrangian of a potentially infinite trajectory rather than an undiscounted $T$-step trajectory. Third, in the following, we apply it not only to the traditional Lagrangian but also to the augmented Lagrangian and the augmented regularised Lagrangian.
@@
Following Theorem 5.5 in Wang et al. (2024), it follows for any value function with cost in $[0,1]$ that
@@
Reformulating the Lagrangian as a value function (see Eq. 5), considering the bounds on the absolute value of the constraint-cost Eq. 13, and observing that the bounds on the (un-)augmented Lagrangian lie in $F_{\lambda},F_{\lambda}]$ (see Eq. 14 and  15), the result can be scaled to obtain
@@
$\mathbf{V}_{\pi,p^{*}}(\rho;\lambda)-\mathbf{V}_{\pi,p^{t}}(\rho;\lambda)$ $\leq\frac{2F_{\lambda}}{t}\left(\frac{M+1}{(1-\gamma)^{2}}\right)\leq\epsilon^{\prime}\,.$
@@ 5.3 Analysis of Approximate TMA
Definition 2 (Action next-state value function (Definition 3.2 in Li et al. (2023))).
@@
It can also be written as $G_{\pi,p}(s,a,s^{\prime})=c(s,a,s^{\prime})+\gamma V_{\pi,p}(s^{\prime})$ (Wang et al., 2024).
@@
The proof follows directly by applying Lemma 5.1 of Wang et al. (2024) (or Lemma 3.4 of Li et al. (2023) with direct parametrisation) to the approximate value function. ∎
@@
$p^{t+1}(\cdot|s,a)=\mathrm{arg}\sup_{p\in\mathcal{P}}\eta_{p}(t)\frac{1}{1-\gamma}d_{\rho}^{\pi,p^{t}}(s,a)\hat{G}_{\pi,p^{t}}(s,a,\cdot)-\alpha_{p}B_{d_{\rho}^{\pi,p^{t}}}(p,p^{t})\quad\text{for all }\mathcal{S}\times\mathcal{A}\,.$ (29)
As a consequence of this update rule, the value difference in $p^{t+1}$ and $p^{t}$ can be bounded. In particular, the ascent property of TMA (Lemma 5.4 of Wang et al. (2024)) is a key result which provides a proof based on rectangularity (Definition 1) to show that TMA increases the objective over iterations. For our purposes, it is used not only in the oracle-based LTMA but also to prove a related ascent property for the Approximate TMA algorithm. The property, which is given below, notes that conditioning on the state or state-action, a pushback property can be achieved, and this can be used to show that (under perfect gradient information) TMA updates always result in increasing value.
Lemma 4 (Ascent property of TMA (Lemma 5.4 of Wang et al. (2024))).
@@
$\langle G_{\pi,p^{t}}(s,\cdot,\cdot),p-p^{t+1}(\cdot|s,\cdot)\rangle+B(p^{t+1}(\cdot|s,\cdot),p^{t}(\cdot|s,\cdot))\leq B(p,p^{t}(\cdot|s,\cdot))-B(p,p^{t+1}(\cdot|s,\cdot))\,$, (30)
@@
$V_{\pi,p^{t}}(\rho)-V_{\pi,p^{t+1}}(\rho)\leq\frac{2\epsilon^{\prime}}{1-\gamma}\,.$ (33)
@@
$\leq\frac{2\epsilon^{\prime}}{1-\gamma}\,.$ (Eq. 32)
@@
Let $V:\mathcal{S}\to\mathbb{R}$ be a value function with instantaneous cost bounded in [C,C][-C,C], fix $\pi\in\Pi$, and let $p^{*}\in\mathrm{arg}\sup_{p\in\mathcal{P}}V_{\pi,p}(\rho)$. Suppose we have an estimate $\hat{G}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ of the action next-state value function which is $\epsilon^{\prime}$-precise as in Eq. 32, and the step sizes follow $\eta_{p}(0)\geq\alpha_{p}(0)\frac{\gamma}{C(1-\gamma)}B_{d_{\rho}^{\pi,p^{0}}}(p^{*},p^{0})$, $\eta_{p}(t)\geq\eta_{p}(t-1)/\gamma$ and $\alpha_{p}(t)=\alpha_{p}(0)>0$ for all $t\geq 1$. It follows that approximate TMA (Algorithm 1) yields a regret bounded by
@@
12:   $p^{t+1}(\cdot|s,a)=\mathrm{arg}\sup_{p\in\mathcal{P}}\eta_{p}(t)\left\langle p,\hat{G}_{\pi,p^{t}}(s,a,\cdot)\right\rangle-\alpha_{p}B_{d_{\rho}^{\pi,p^{t}}}(p,p^{t})$ for all $\mathcal{S}\times\mathcal{A}$.
@@ 5.4 Analysis of Robust Sample-based PMD-PD
Having shown results Lagrangian TMA in oracle-based and sample-based settings, we now turn to proving the sample complexity of Robust Sample-based PMD-PD (Algorithm 3). The algorithm modifies Sample-based PMD-PD (Liu et al., 2021b) by accounting for the mismatch between the nominal transition kernel and the optimal transition kernel in the min-max problem in Eq. 4. The complexity analysis below is slightly simplified by dropping purely problem-dependent constants (e.g. $\gamma$, $\eta$, $\zeta$) from the big-O statements.
@@
As sample-based algorithms work with approximate value functions, Lemma 9 describes the conditions under which one can obtain an $\epsilon$-approximation of the cost functions and the regularised augmented Lagrangian. The same settings as sample-based PMD in Lemma 15 of Liu et al. (2021b) transfer to the Robust Sample-based PMD algorithm. We rephrase the lemma to include updated transition dynamics.
Lemma 9 (Value function approximation, Lemma 15 of Liu et al. (2021b) rephrased).
@@
Lemma 10 (Bound on regularised augmented Lagrangian under approximate entropy-regularised NPG, Lemma 18 in Liu et al. (2021b)).
@@
The proof makes use of Lemma 20. In particular, it notes that $K=\Theta(1/\epsilon)$ and the value difference to $\pi_{k}^{*}$ is bounded by $1/K$ after $t_{k}$ steps. After applying the pushback property (Lemma 17) to $\pi$, $\pi_{k}$, and $\pi_{k}^{*}$, and derives an upper bound on the Bregman divergence w.r.t. $\pi_{k}^{*}$ based on $B_{d_{\rho}^{\pi,p_{k}}}(\pi,\pi_{k+1})+\left\lVert\log(\pi_{k}^{*})-\log(\pi_{k}^{t+1})\right\rVert_{\infty}$, where the second term is $\Theta(1/K)$. We refer to Lemma 18 of Liu et al. (2021b) for the detailed proof. ∎
@@
Choose parameter settings according to Lemma 9, let $\alpha=\frac{2\gamma^{2}m\eta_{\lambda}}{(1-\gamma)^{3}}$, $\eta_{\lambda}=1$, and $\eta=\frac{1-\gamma}{\alpha}$. Moreover, let $\pi^{*}\in\mathrm{arg}\inf_{\pi\in\Pi}\Phi(\pi)$, let $\epsilon_{k}^{\prime}>0$ be the upper bound on the LTMA error tolerance at macro-iteration $k$ such that $\epsilon_{k}^{\prime}=\Theta(\epsilon)$, and let $\mathcal{P}$ be a rectangular uncertainty set according to Lemma 24a) or b). Then within a total number of queries to the generative model equal to $T=\tilde{\mathcal{O}}(\epsilon^{-3})$, Robust Sample-based PMD-PD provides the following guarantees:
(a) average regret upper bound:
@@
(b) average constraint-cost upper bound:
@@
c) robust Lagrangian average regret upper bound under (oracle-based) LTMA: under the conditions of Lemma 2, with a number of LTMA iterations $t_{k}^{\prime}=\Theta\left(\frac{F_{\lambda_{k}}M}{\epsilon_{k}^{\prime}(1-\gamma}\right)$, and with $\mathcal{P}$ defined according to the preconditions of Lemma 24b), we have
@@
d) robust Lagrangian average regret upper bound under Approximate LTMA: under the conditions of Lemma 7 as applied to the Lagrangian at each macro-iteration, with $C=F_{\lambda}$, and with $\mathcal{P}$ defined according to the preconditions of Lemma 24b), we have
@@
(a) The proof uses Lemma 10 and Lemma 15 to obtain a telescoping sum similar to the analysis of Theorem 3 in Liu et al. (2021a). To account for transition kernel changes, it then upper bounds the cumulative regret difference using Lemma 24) and the Bregman divergence terms in the RHS of Eq. 39 using Lemma 25. This then results in a telescoping sum.
@@
Sample complexity: To obtain the sample complexity under Algorithm 3 with Approximate TMA, plug in the settings from Lemma 9 and Lemma 7 to obtain
$T=\sum_{k=1}^{K}\left(M_{V,k}N_{V,k}+\sum_{t=0}^{t_{k}-1}M_{Q,k}N_{Q,k}+\sum_{t=0}^{t_{k}^{\prime}-1}M_{G,t}N_{G,t}\right)=\tilde{\mathcal{O}}(\epsilon^{-3})\,.$ (42)
5.5 Extension to non-rectangular uncertainty sets
Similar to arguments made by Xiao (2022) for Policy Mirror Descent, oracle-based TMA has a direct relation to the policy iteration algorithm.
Lemma 11 (Relation between TMA and Policy Iteration).
Let $\eta_{p}(t)\to\infty$ for all $t\in\{0,\dots,t_{k}^{\prime}-1\}$. Then under the direct parametrisation, TMA is equivalent to Policy Iteration as defined for transition kernels in Li et al. (2023).
Proof.
Note that for $\eta_{p}(t)\to\infty$, applying the TMA update rule, analogous to Eq. 29,
$p^{t+1}(\cdot|s,a)$ $=\mathrm{arg}\sup_{p\in\mathcal{P}}\eta_{p}(t)\frac{1}{1-\gamma}d_{\rho}^{\pi,p^{t}}(s,a)G_{\pi,p^{t}}(s,a,\cdot)-\alpha_{p}B_{d_{\rho}^{\pi,p^{t}}}(p,p^{t})$
$=\mathrm{arg}\sup_{p\in\mathcal{P}}\left\langle p,G_{\pi,p^{t}}(s,a,\cdot)\right\rangle-\frac{\alpha_{p}}{\eta_{p}(t)}B(p,p^{t})$ (reweighting similar to Eq. 4)
$=\mathrm{arg}\sup_{p\in\mathcal{P}}\left\langle p,G_{\pi,p^{t}}(s,a,\cdot)\right\rangle\,$,
which is equivalent to policy iteration. The latter can also be written as $\mathrm{arg}\sup_{p\in\mathcal{P}}\left\langle p-p^{t},G_{\pi,p^{t}}(s,a,\cdot)\right\rangle$, which based on Lemma 3 is equivalent to
$\mathrm{arg}\sup_{p\in\mathcal{P}}\left\langle p-p^{t},\nabla V_{\pi,p^{t}}(\rho)\right\rangle$ (43)
as in Eq. 3.4 of Li et al. (2023). ∎
Algorithm 2 Conservative Policy Iteration over transition kernels using Approximate TMA
1:Inputs: $\epsilon^{\prime}>0$, policy $\pi\in\Pi$, value function $V:\mathcal{S}\to\mathbb{R}$, initial transition kernel $p^{0}$, compact and convex uncertainty set $\mathcal{P}$, learning rate $\eta_{p}(t)\to\infty$ for all $t\geq 0$.
2:while True do
3:  One step of Approximate TMA with learning rate $\eta_{p}(t)$ to obtain $p_{\epsilon^{\prime}}$ as $\epsilon^{\prime}$-optimal solution to Eq. 43.
4:  Compute Frank-Wolfe gap $\mathbb{G}_{t}=\langle\nabla V_{\pi,p^{t}}(\rho),p_{\epsilon^{\prime}}-p^{t}\rangle$.
5:  if $\mathbb{G}_{t}\leq\epsilon^{\prime}$ then
6:   return $\hat{p}=p^{t}$.
7:  else
8:   Let $\beta_{t}=\mathbb{G}_{t}\frac{(1-\gamma)^{3}}{4\gamma^{2}}$
9:   Let $p^{t+1}=(1-\beta_{t})p^{t}+\beta_{t}p_{\epsilon^{\prime}}$.
10:   $t\leftarrow t+1$.
11:  end if
12:end while
Exploiting the above relation, and using Conservative Policy Iteration to transition kernels according to Algorithm 2, we can obtain the sample complexity under non-rectangular RCMDPs using arguments by Li et al. (2023). In particular, while Wang et al. (2024) provide an analysis of TMA only for $s$-rectangular RMDPs and derives oracle-based convergence rates for the policy optimiser, it is possible to extend the analysis to non-rectangular uncertainty set by relating it to the smallest $s$-rectangular uncertainty set that contains it. With this reasoning, the result below considers the Appproximate TMA algorithm for non-rectangular RCMDPs and then derives a full sample-based analysis of the full algorithm optimising all three parameters $(\pi,p,\lambda)$. The Algorithm 2 requires $t_{k}^{\prime}=\mathcal{O}(\epsilon^{-2})$ iterations of the Approximate TMA algorithm. Therefore, with derivations as in Theorem 2, a sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-5})$ is obtained. In addition to the added complexity, the non-rectangularity also comes with an additional error term based on the degree of non-rectangularity $\delta_{\mathcal{P}}$, defined as
$\delta_{\mathcal{P}}=\max_{p^{\prime}\in\mathcal{P}}\left(\max_{p_{s}\in\mathcal{P}_{s}}\langle\nabla\mathbf{V}_{\pi,p^{\prime}},p_{s}-p^{\prime}\rangle-\max_{p\in\mathcal{P}}\langle\nabla\mathbf{V}_{\pi,p^{\prime}},p-p^{\prime}\rangle\right)\,$, (44)
where $\mathcal{P}_{s}$ is the smallest $s$-rectangular uncertainty set that contains $\mathcal{P}$. In the algorithm, the supremum and maximum are equivalent since $\mathcal{P}$ and $\mathcal{P}_{s}$ are assumed to be compact. Using compactness and convexity of these sets, Remark 3 of Li et al. (2023) shows that $\delta_{\mathcal{P}}$ is finite using Berge’s maximum theorem.
Theorem 3 (Sample complexity for non-rectangular RCMDPs).
Let $\mathcal{P}$ be a non-rectangular uncertainty set according to Lemma 24 a),b) or c). Further, assume the preconditions of Theorem 2 and Lemma 30. Then applying Algorithm 2 to solve the inner problem on the Lagrangian, Robust Sample-based PMD-PD obtains an average regret of
$\frac{1}{K}\sum_{k=1}^{K}\Phi(\pi_{k})-\Phi(\pi^{*})$ $=\mathcal{O}(\epsilon+\delta_{\mathcal{P}})\,$
within $\tilde{\mathcal{O}}(1/\epsilon^{5})$ samples.
Proof.
Let $k\in[K]$. Applying Lemma 30 to $\pi_{k}$, Algorithm 2 obtains a solution $p_{k}=\hat{p}$ such that
$\mathbf{V}_{\pi_{k},p^{*}}(\rho;\lambda_{k-1})-\mathbf{V}_{\pi_{k},p_{k}}(\rho;\lambda_{k-1})\leq M(2\epsilon_{k}^{\prime}+\delta_{\mathcal{P}})=\mathcal{O}(\epsilon+\delta_{\mathcal{P}})$
within $t_{k}^{\prime}=\tilde{\mathcal{O}}(n/\epsilon^{2})$ samples, where $n$ indicates the number of samples to evaluate Eq. 43. Following the same argumentation as in Theorem 2d) and adding the additional term $\mathcal{O}(\delta_{\mathcal{P}})$,